OpenStudy (anonymous):

Professor Dupré Calc gives her class 20 study questions, from which she will select 8 to be on the final exam. How many ways can she select the questions?

OpenStudy (anonymous):

20c8 ways i think so

OpenStudy (anonymous):

yes \[\binom{20}{8}\] sometimes written as \(_{20}C_8\) and read "20 choose 8"

OpenStudy (anonymous):

you compute via \[\frac{20\times 19\times 18\times 17\times 16\times 15\times 14\times 13}{8\times 7\times 6\times 4\times 3\times 2}\] cancel first, multiply last or use a calculator http://www.wolframalpha.com/input/?i=20+choose+8

OpenStudy (anonymous):

think i had a typo above, there should also be a 5 in the denominator